OpenStudy (anonymous):

so if in an equation has 2 unknown values is it still unfactorable?

OpenStudy (anonymous):

you can't solve a single equation with more than one variable (unknown), but that has nothing to do with factoring. you may factor an expression with more than one variable

OpenStudy (amistre64):

x^2 - y^2 can be factored

OpenStudy (anonymous):

for example: X^2+2ax+a^2

OpenStudy (anonymous):

yes, that is factorable as (x+a)^2

OpenStudy (anonymous):

let me try first and if I can't factor it I'll come back. :D

OpenStudy (anonymous):

amistre's example also is factorble into (x-y)(x+y)

OpenStudy (anonymous):

but you can't SOLVE 2x+5y=10, for example

OpenStudy (anonymous):

ok can u tell me how did you factor this equation??thanks

OpenStudy (anonymous):

amistre's example is a difference of two squares; it factors according to a formula\[a^2-b^2=(a-b)(a+b)\]you can check this by multiplying the RHS (FOIL) and you will get the LHS.

OpenStudy (anonymous):

your example is a perfect square trinomial\[(p+q)^2=p^2+2pq+q^2\]

OpenStudy (anonymous):

in your example p=x and q=a

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

np:})